OpenStudy (abbycross167):

Can someone please check the answer to my question? Please? The length of a rectangle is (2x - 7) inches and its width is (x - 5) inches. Find an expression for the perimeter of the rectangle. a. 3x - 12 b. 6x + 70 c. 3x + 35 d. 6x - 24

OpenStudy (abbycross167):

@AloneS. @jhonyy9 could one you y'all help me?

OpenStudy (abbycross167):

I think it is B or D

jhonyy9 (jhonyy9):

are you here please ?

OpenStudy (abbycross167):

yes sir/ma'am, I have been waiting to see what you think of it

jhonyy9 (jhonyy9):

i m man - so what is the formula for perimeter of a rectangle ?

OpenStudy (abbycross167):

P=2(l+w)

jhonyy9 (jhonyy9):

how you get the perimeter - ok - yes right so than using these above given dates what will be P= ?

jhonyy9 (jhonyy9):

l=2x-7 w=x-5 P=?

jhonyy9 (jhonyy9):

use formula what you have wrote above please

OpenStudy (abbycross167):

how would I do that with the variable x?

jhonyy9 (jhonyy9):

indifferent this is just an expression of the side length and width length - ok. ?

jhonyy9 (jhonyy9):

you need just substitute the given dates inside formula what you have wrote above for perimeter - do you can substituting the value of the length in place of l and the value of the width in place of w - ok. ?

jhonyy9 (jhonyy9):

P=2(l+w) so please substitute the given dates in place of l and w - so what will get ?

OpenStudy (abbycross167):

P=2((2x - 7) + (x - 5))

OpenStudy (jdoe0001):



OpenStudy (abbycross167):

@jdoe0001 thank you very much for making this! Could you please try to help me answer this question? I don't know how

OpenStudy (jdoe0001):

just add them up :)

OpenStudy (abbycross167):

oh, I tried that but I got -28 for 2x-7 (which is -14 (right?)) so it'll be -28 for both sides

OpenStudy (abbycross167):

@jdoe0001 could you help me??